# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

### 05 Feb 2024 v1

@@ 4.1.1. DG1: Contextual cues
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). Understanding the paper and critically reflecting on the content are essential for decision-making and review writing in the later stage. Previous research showed that providing questions can increase user engagement in actively searching for information. Previous systems mostly use template-based guided questions to engage readers (Peng et al., 2022). To Offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper [DG1], ReviewFlow provides two types of contextual cues for novices, as shown in Figure 3. One type is section level reflection cues guided by community criteria which are adapted based on each paper section’s content together with the review criteria. For section-level reflection cues in the Figure 7-A), it took the entire section, the paper abstract, and the community criteria into account to generate questions. Readers can either read these questions before they read the section to obtain contextual guidance or after to reflect on the section’s content. Another type is phrase level cues adapted to paper content. As shown in Figure 7-B, participants have the freedom to highlight the content that they would like to reflect on at the phrase level and then select the review criteria. Then, it generated cues for participants to reflect on.
Figure 3. Contextual cues. ReviewFlow provides (A) section level cues guided by community criteria and (B)phrase level cues adapted to paper content and selected criteria
@@ 4.1.2. DG2: In-situ citation recommendation as knowledge scaffolding
Previous research in scientific literature reviewing highlights the importance of in-situ knowledge support (August et al., 2022; Kang et al., 2022; Rachatasumrit et al., 2022; Chang et al., 2023). For example, PaperPlain provided in-situ plain language section summaries to facilitate paper reading. To enable in-situ knowledge support for assessing novelty compared to prior work, ReviewFlow in-situ presents TLDR summary when the user click on each reference to provide background knowledge around each paper, as shown in Figure 4-D. The TLDR is queried from semantic scholar (sem, 2015). To further raise the awareness of the existing reference in the reading workflow, ReviewFlow provided a popup window with potential missing citations from the same venues, as shown in Figure 4-C.
Figure 4. In-situ knowledge scaffolding. ReviewFlow provides: (C) a list of relevant paper from the same venue that are not cited and (D) in-situ citation summary including title, author and TLDR
@@ 4.1.3. DG3: Notes-to-outline synthesis
In the review process, we observe expert reviewers synthesize notes to make a plan on how to draft a review. To facilitate the iterative planning and revising process in writing, intelligent systems further use the chain of thoughts prompting method to break down the large problem into step-by-step prompts (Wei et al., 2022). Specifically, previous research used the chain of thought approach to decompose the writing tasks where they first generate an outline and then automatically turn the outline into the story generation. The evaluation demonstrated that this decomposition approach can improve the coherence of long story generation (Yang et al., 2022b, a). However, this approach could be highly human controllable where humans can control the story generation by modifying the outlines. Recent studies explored the potential of automatically generating feedback using LLM for conference papers (Liang et al., 2023).
Figure 5. Scaffolding features support the review drafting process. The top part shows the Flower and Hayes cognitive process of writing (Flower and Hayes, 1981). The bottom part shows the ReviewFlow features that support each stage of writing. On the left, ReviweFlow summarized notes into broad topics under strengths and weaknesses to help reviewers plan the writing. In the middle, the topics can expand to a detailed outline. On the right, it encourages self-reflection and allows post-editing
By modeling the expert workflow of reading, note-taking, and synthesizing practiced by experienced reviewers, we designed the notes-to-outline synthesis feature that facilitates notes synthesis and review planing (Flower and Hayes, 1981). Figure 5 showed that ReviewFlow first synthesized notes to high-level outline to help reviewers organize their notes into structured topics and provide ideas for reviewers to make decisions. To further guide the reviewing process such that the work aligns with community standards, notes are summarized into broad topics and structured according to community standards, which includes aspects of summary, strengths, and weaknesses. Given these high-level outlines, reviewers can add their own points or delete any points. If reviewers would like to expand the high-level outline into a detailed outline with complete bullet points, reviewers can click the expand button that synthesized notes into a structured outline, as shown in Figure 5.
@@ 4.1.4. DG4: Fact-checking between outline and source notes and self-reflections
When the research team introduced the idea of using AI to facilitate the reading and writing process, participants expressed the need for fact-checking and providing explanations. To avoid biasing the decision to either accept or reject the paper, but encourage justifications, ReviewFlow provided the feature that for each synthesized outline bullet, when the reviewers hover on it, corresponding notes used to generate the topic will be highlighted in the middle column. The pdf on the left column will also auto-scroll to shows the location of the notes, as shown in Figure 6.
Figure 6. When the participants hover on each outline element, ReviewFlow presents the visual mapping between the summarized outline and source notes. This visual mapping facilitate fact-checking and comparison between summarised topics and original notes.
To help participants check whether the tone of the review is appropriate as well as the structure of the written review meet the community standard, ReviewFlow provided self-reflection features right before the participants submitted their review. As shown in the right column of the Figure 5, before participants submit their final reviews, they need to self-reflect on each aspect listed on the reflection card. Participants can go back and revise their draft if they are not satisfied with the review.
@@ 4.2. System implementation
ReviewFlow’s front-end is a React web application responsible for displaying the PDF, managing the annotation data, and displaying the text input used to write the review draft. The back-end uses a Flask server which handles the GPT-4 API endpoints, MongoDB endpoints for storing data, and GROBID for PDF data extraction. This data includes all of the parsed text of the PDF, its PDF coordinates, and citations linked to their references.
@@ 4.2.1. Contextual cues
After the user uploads a paper PDF on the front-end, a request is sent to ReviewFlow’s back-end server where GROBID is used to parse and extract the paper’s content, creating an XML/TEI structured document with the coordinates and content of the section titles, text body sentences, and inline citations. This extracted data is then sent to the front-end client where ReviewFlow combines GROBID’s section data with PDF.js’s section data to make GPT-4 API calls to generate the contextual reflection questions for each of the paper’s sections.
To generate phase-level cues, when the user highlights text on the PDF and clicks on the button with the light bulb icon, a pop-up appears with the 4 critical review aspects: importance, novelty, validity, and clarity. When the user selects one of the aspects and clicks “Get a Question”, GPT-4 will use the selected aspect, highlighted phases, paragraph of the highlighted phases, and paper abstract to generate cues.
@@ 4.2.2. In-situ citation recommendation
When the PDF is initialized, ReviewFlow not only utilizes GROBID’s XML/TEI data to create markers for the section-specific reflection questions but also uses the data to create a citation layer that overlays all in-line citations. When the user hover over each citation, a popup shows the paper title, publication date, DOI link and the tldr description that is searched using Semantic Scholar API (sem, 2015). To recommend corresponding papers, the system calls the Semantic Scholar Recommendation API [^2] https://api.semanticscholar.org/api-docs/recommendations and uses the keywords of the paper together with the venue to retrieve the most similar paper. After checking the retrieved papers are not included in the current paper, the top three papers are added as a citation pop-up.
@@ 4.2.3. Notes-to-outline Synthesis
After the user creates multiple notes, a “Summarize Notes” button will appear at the bottom of the notes panel. Upon clicking this button, the front-end client will organize all the user’s notes data including the corresponding highlighted paper content, selected tags, note text, and paper abstract. Subsequently, the system server will use this data to prompt the GPT-4 model. We used the few-shot learning paradigm with the corresponding prompts (Liu et al., 2023b): ”Please create three important topics on the paper’s [strengths] and [weaknesses] with less than ten words that combine and summarize the user notes.” This output is then formatted in JSON and streamed directly into the front-end’s text-editable draft panel with a summary and topics organized on strengths and weaknesses. When the user clicks on the ”Expand” button at the bottom of the draft panel, all of the text in the draft text input box, notes, and paper abstract is sent to the back-end server where a GPT-4 API call is requested. The GPT-4 API call will then respond with a streamed JSON-formatted output with more details based on the user’s notes and topics specifically for the Strength and Weakness sections.
@@ 5.1. Study Design
We conducted a within-subjects experiment with 16 participants where each participant will experience both conditions in two sessions separately. We counterbalanced the order of two conditions and papers using Latin Square to minimize the potential order effect. To reduce knowledge transfer and minimize fatigue, we scheduled the two study sessions of each participant at least 24 hours apart.
@@ 5.1.1. Participants
Before the study, we sent out a pre-study survey to participants to collect information about their expertise, research topics, review experience, ML knowledge level, and demographics (e.g. age, gender, race). We recruited 16 participants who have experience in conducting academic research for at least two years and have zero to two years of conference peer review experience. We advertised recruitment messages to colleagues, mailing lists, communication channels, and social media and recruited participants from four universities across the US. Using a snowball sampling approach, we asked participants to refer their friends and colleagues. Participants have on average 1.2 years of writing and submitting academic papers. All participants were compensated for \$20 per hour. The study is IRB-approved.
@@ 5.1.2. Procedure
Before the user study, participants filled out a pre-survey that captured their previous experience in reviewing and reviewing papers on ML conferences and their expertise in ML to join the study through an online teleconferencing tool. We asked participants whether they had read the paper before to make sure they all were seeing the work for the first time. In the beginning, participants read and virtually signed our IRB-approved consent form. Then they conducted reviews on two papers using different versions of our interface (ReviewFlow or Baseline). To counterbalance the order effect, we randomized the order of the control condition and the experiment condition for each participant, so half participants encountered ReviewFlow in their first session, and the other half experienced it in their second session. For each session, we followed the review process used at the CSCW conference, where we provided the paper draft and all three individual reviews with their ratings and confidence scores.
For both sessions, participants were told to spend around 60 minutes – and no less than 30 minutes – on the review, but they could take as much time as needed. In the pilot study with 4 participants who had reviewing experience before the study, we found that participants can finish writing reviewing within 45 minutes. Before each condition session, we gave the participants a quick 2-minute demo of the interface, while in the experiment session demo, we also introduced the functionality of ReviewFlow. After two sessions, the research team asked the participants to fill out a post-survey to evaluate the system, such as their satisfaction with each feature and their trust in the system. The research team then conducted a 15-minute semi-structured post-interview to ask open-ended questions about their experience, perceptions, and feedback. The post-interviews were video recorded with participants’ permission and were transcribed into text for later analysis. Each session takes around 75 minutes.
@@ 5.1.3. Paper Selection Process
We collected recent one-year papers from HCI-related conferences including CHI, CSCW, UIST, UBICOMP, IUI, DIS using the list provided by  [^3] http://www.conferenceranks.com/. To make our study time short so that people have the energy to finish the task, we filtered papers that had fewer than 3000 words. Then we filtered out papers that have technical terms and jargon. The two papers we selected need to have similar lengths, similar difficulties and from the same conference. Combining all the criteria above, we selected two comparable papers from CSCW Companion that were used for the within-subject experiment. We measured participants’ existing knowledge of the two papers as a control variable. During the study, participants described their knowledge of each paper’s topic from not familiar as 1 to very familiar as 7. All participants reported that they had never read or remembered the papers before.
@@ 5.1.4. Measures
We collected a mix of quantitative and qualitative data, including each participant’s log data that captured their interactive behaviors with the system, the final review artifacts for each paper (N=32) that recorded the final review, the post-survey (N=32 and the interview transcripts (N=16). The research team analyzed these combined sources of data to reveal insights towards our research questions.
@@
To measure the quality of the review, we recruited two experts who have conducted research for more than three years of review experience in HCI/CSCW conferences. They counted the number of strengths and weakness in the review (a proxy for coverage) and rated the quality of all final reviews (N=32) with a five-dimension rubric based on reviewer guidelines for CSCW conference and previous research (Yuan et al., 2021):
@@
Two experts also rated these five dimensions on a simple seven-point scale (1-7). Each expert first read one paper and wrote a peer review of the paper. After this process, the research team first provided two examples and provided instructions for them to rate and discuss until they reached a consensus on ratings based on the instructions. Then they rate each dimension of the review on the paper they wrote independently (N=16). The inter-rater reliability between two experts on all 32 data is moderate where Krippendorff’s alpha are higher than 0.50 (Krippendorff, 2018). The research team then used the average scores by two experts for each paper dimension.
@@
To measure participants’ interaction with the tool, we instrumented the interface to collect a range of user activity log data. We collected two timing measures – how long each participant took to finish the review session and how long each participant spent editing the review within the text box. The entire review session time includes reading time and writing time. The ReviewFlow interface also collected interaction data to indicate how much each participant interacted with each scaffolding feature, including how many times they answered the contextual cues, clicked on the citation summary, checked the citation pop-ups and used the outline summarization feature.
@@
To evaluate users’ preferences for the ReviewFlow experience compared to a plain review Editor, we asked participants to fill out a short post-study survey. The survey asked participants to directly compare the perceived usefulness, enjoyment, easiness, and sense of control between the ReviewFlow and baseline system. Then we specially asked participants to evaluate each of the machine-generated features. The survey collected their 7-point Likert scale ratings on perceived usefulness, and perceived accuracy for each feature. We collected the level of cognitive demand using NASA TLX on a scale of 1-7 (Hart and Staveland, 1988). We further collected the feeling of control and collaboration on a scale of 1-7.
Previous research showed that scaffolding can promote learners’ self-efficacy (Yantraprakorn et al., 2013). Here, we further measure whether using the scaffold system can improve novice reviewers’ self-efficacy and knowledge. We asked each participant to report self-efficacy after using the ReviewFlow system and the Baseline system by answering the question “How confident are you in your ability to write a conference peer review next time after using the system( 1 as not confident and 7 as very confident)”.
After the post-survey, we conducted a 15-minute semi-structured interview with all participants to capture their overall thoughts as well as specific perceptions of machine-generated highlights and summaries. For example, the research team asked “What do you think of the difference between the task with and without the support of ReviewFlow”, and “Which function did you use the most? Which one did you like the most?”, and “What concerns did you have when using ReviewFlow?”.

### 26 Feb 2024 v2

@@ 4.1.1. DG1: Contextual cues
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). Understanding the paper and critically reflecting on the content is essential for decision-making and review writing in the later stage. Prior research showed that providing questions can increase user engagement in actively searching for information. Prior systems mostly used template-based guided questions to engage readers (Peng et al., 2022). To offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper [DG1], ReviewFlow provides two types of contextual cues for novices, as shown in Figure 3. One type is section-level cues guided by community criteria which are adapted based on each paper section’s content together with the review criteria. For section-level reflection cues in Figure 3-A), it takes the entire section, the paper abstract, and the community criteria into account to generate questions. Users can either read these questions before they read the section to obtain contextual guidance or after to reflect on the section’s content. Another type is phrase-level cues adapted to paper content. As shown in Figure 3-B, users have the freedom to highlight the content that they would like to reflect on at the phrase level and then select the review criteria. Then, it generates cues for participants to reflect on in real-time.
Figure 3. Contextual cues. ReviewFlow provides (A) section-level cues guided by community criteria and (B) phrase-level cues adapted to user highlighted content and selected criteria.
@@ 4.1.2. DG2: In-situ citation recommendation as knowledge scaffolding
Prior research in scientific literature review highlights the importance of in-situ knowledge support (August et al., 2022; Kang et al., 2022; Rachatasumrit et al., 2022; Chang et al., 2023). To enable in-situ knowledge scaffolding for assessing novelty compared to prior work, when the user clicks on each reference, ReviewFlow in-situ presents the title and a TLDR summary, as shown in Figure 4-D. The TLDR is queried using Semantic Scholar (sem, 2015). To raise awareness of the existing references in the reading workflow, ReviewFlow provides a popup window with potential missing citations from the same venues, as shown in Figure 4-C.
Figure 4. In-situ knowledge scaffolding. ReviewFlow provides (C) a list of relevant papers from the same venue that are not cited and (D) an in-situ citation summary including title, author, and the TLDR summary.
@@ 4.1.3. DG3: Notes-to-outline synthesis
In the review process, we observed expert reviewers synthesize notes to make a plan on how to draft a review. By modeling the cognitive processes of writing proposed by Flower and Hayes, we designed the notes-to-outline synthesis feature (Flower and Hayes, 1981). As shown in Figure 5, we decomposed the cognitive processes which include planning, translating, and reviewing, and provided scaffolding features for each stage. In the planning stage, to facilitate writers to generate ideas, set goals, and organize information, ReviewFlow provides a notes-to-topic synthesis that summarizes the user notes together with the highlighted text into general topics, such as “needs more detailed citation description”. These topics are organized into a structure the aligned with the community practices including summary, strength, and weakness. In the translating stage, to facilitate drafting, ReviewFlow can expand broad topics into detailed bullets if the user clicks the expand button. These bullet points provided a more specific summary based on users’ notes. In the reviewing stage, to help users evaluate their written review and revise the text, ReviewFlow pops up a self-reflection card and asks users to self-reflect on their written review using the review criteria, including tone, comprehensive, constructive, justified, and accurate. This encourages users to revise the text if they would like to improve the review quality.
Figure 5. Scaffolding features support the review writing process. The top part shows Flower and Hayes cognitive process of writing (Flower and Hayes, 1981). The bottom part shows the ReviewFlow features that support each writing step. On the left, ReviewFlow summarized notes into broad topics under strengths and weaknesses to facilitate planning. In the middle, use can click to expand the topics to a detailed outline. On the right, the pop up window encourages self-reflection and post-editing.
@@ 4.1.4. DG4: Fact-checking between outline and source notes and self-reflections
When the research team introduced the idea of using AI to facilitate the reading and writing process, participants expressed the need for fact-checking and providing explanations. To avoid biasing the decision to either accept or reject the paper but encourage justifications, ReviewFlow provided the feature that for each synthesized outline bullet, when the reviewers click it, the notes in the middle column will be highlighted. The PDF on the left column will also scroll automatically and highlight the corresponding location of the notes, as shown in Figure 6.
Figure 6. When the participants click on each outline element, ReviewFlow shows the visual mapping among the summarized outline bullet (a), the note in the middle (b) and original pdf with highlights (c).
@@ 4.2. System implementation
ReviewFlow’s front-end is a React web application built on top of an existing PDF highlighting library  [^2] https://github.com/agentcooper/react-pdf-highlighter. The front end is responsible for displaying the PDF, managing the annotation data, and displaying the text input used to write the review draft. The back-end uses a Flask server which handles the GPT-4  [^3] https://openai.com/research/gpt-4 API endpoints, MongoDB endpoints for storing data, and GROBID (gro, 2008) for PDF data extraction. This data includes all of the parsed text of the PDF, its PDF coordinates, and citations linked to their references. Detailed prompts are provided in the Appendix.
@@ 4.2.1. Contextual cues
After the user uploads a paper PDF on the front end, a request is sent to ReviewFlow’s back-end server where GROBID is used to parse and extract the paper’s content, creating an XML/TEI structured document with the coordinates and content of the section titles, text body sentences, and inline citations. This extracted data is then sent to the front-end client where ReviewFlow combines GROBID’s section data with PDF.js’s section data to make GPT-4 API calls to generate the contextual cues for each of the paper’s sections.
Each GPT-4 API call utilizes the section’s text as a prompt to generate a contextual reflection question for each of the following critical review aspects: importance, novelty, validity, and clarity. The GPT-4 response is formatted as a JSON and is streamed to the front-end client. To generate phrase-level cues, when the user highlights text on the PDF and clicks on the button with the light bulb icon, a pop-up appears with the review criteria. When the user selects one of the aspects and clicks ”Get a Question”, GPT-4 will use the selected aspect, highlighted phrases, paragraph of the highlighted phrases, and paper abstract to generate cues.
@@ 4.2.2. In-situ citation recommendation
When the PDF is initialized, ReviewFlow uses GROBID’s XML/TEI data to create a citation layer that overlays all in-line citations. When the user clicks on an in-text citation, a popup shows the paper title, publication date, DOI link, and a short description that is searched using Semantic Scholar API (sem, 2015). To recommend corresponding papers, the system calls the Semantic Scholar Recommendation API [^4] https://api.semanticscholar.org/api-docs/recommendations and uses the keywords of the paper together with the venue to retrieve the most similar paper. After filtering out the retrieved papers that have already been included in the current paper, the top three papers are added as a citation pop-up.
@@ 4.2.3. Notes-to-outline Synthesis
After the user creates multiple notes, a “Summarize Notes” button will appear at the bottom of the notes panel. Upon clicking this button, the front-end client will organize all the user’s notes data including the corresponding highlighted paper content, selected tags, note text, and paper abstract. Subsequently, the system will use this data as prompt. We use the few-shot learning paradigm with the corresponding prompts (Liu et al., 2023b): “Please create three important topics on the paper’s [strengths] and [weaknesses] with less than ten words that combine and summarize the user notes.” The output is then formatted in JSON and streamed directly into the front-end’s text-editable draft panel with a summary and topics organized on strengths and weaknesses. When the user clicks on the “Expand” button at the bottom of the draft panel, all of the text in the draft text input box, notes, and paper abstract is sent to the back-end server and requested a GPT-4 API call. The GPT-4 API call will then respond with a streamed JSON-formatted output with more details based on the user’s notes and topics specifically for the strength and weakness sections.
@@ 5.1. Study Design
We conducted a within-subjects experiment with 16 novice participants where each participant experienced both the ReviewFlow condition and the Baseline condition in two sessions separately. We counterbalanced the order of two conditions and papers using a Latin Square design to minimize the potential order effects. To reduce knowledge transfer and minimize fatigue, we scheduled the two study sessions of each participant at least 24 hours apart. The ReviewFlow condition includes all scaffolding features, while the Baseline condition includes the minimal guideline. In the baseline interface, users can still highlight, take notes, and tag notes. Community guideline is also provided.
@@ 5.1.1. Participants
Before the study, we sent out a pre-study survey to participants to collect information about their expertise, research topics, review experience, knowledge level on AI, and demographics (e.g. age, gender, race). We recruited 16 participants who have experience in conducting academic research for at least two years and have zero to twice conference peer review experience to make sure that they are novice reviewers who lack review experience but are equipped with domain knowledge. We advertised recruitment messages to colleagues, mailing lists, communication channels, and social media and recruited participants from four universities across the US. Using a snowball sampling approach, we asked participants to refer their friends and colleagues. Participants have on average 1.2 years of writing and submitting academic papers. Each study section takes around 90 minutes and all participants were compensated for \$20 per hour. The study is IRB-approved.
@@ 5.1.2. Procedure
Before the user study, participants filled out a pre-survey that captured their previous experience in reviewing and reviewing papers on HCI/CSCW conferences and their expertise in HCI/CSCW. We asked participants whether they had read the paper before to make sure they all were seeing the work for the first time. Participants conducted reviews on two papers using different versions of our interface (ReviewFlow or Baseline). To counterbalance the order effect, we randomized the order of the control condition and the experiment condition for each participant, so half participants encountered ReviewFlow in their first session, and the other half experienced it in their second session. For each session, we followed the review process used at the CSCW conference, where we provided the paper draft and the review guidelines.
For both sessions, participants were told to spend around 60 minutes – and no less than 30 minutes – on the review, but they could take as much time as needed. In the pilot study with 2 participants who had very little review experience, we found that participants could finish reviewing within 45 minutes. Before each condition session, we gave the participants a quick 2-minute demo of the interface, while in the experiment session demo, we also introduced the functionality of ReviewFlow. After each session, the research team asked the participants to fill out a post-survey to evaluate the system and assess their self-efficacy. After the second session, we conducted a 15-minute semi-structured post-interview to ask open-ended questions about their experience, perceptions, and feedback. The post-interviews were video recorded with participants’ permission and were transcribed into text for later analysis.
@@ 5.1.3. Paper Selection Process
We collected recent one-year papers from HCI-related conferences including CHI, CSCW, UIST, UBICOMP, IUI, DIS using the list provided by  [^5] http://www.conferenceranks.com/. To make our study time short so that people have the energy to finish the task, we filtered papers that had fewer than 3000 words and further filtered out papers that had technical terms and jargon. The two papers we selected need to have similar lengths, similar difficulties, and from the same conference venues. Combining all the criteria above, two papers from CSCW Companion were selected for the study. The first paper includes the keywords “virtual studying, video streaming, awareness” and contains 2853 words. The second paper includes the keywords “virtual environment, cross-lingual collaboration, team formation” and contains 2910 words.
@@ 5.1.4. Measures
We collected a mix of quantitative and qualitative data, including each participant’s log data that captured their interactive behaviors with the system, the final review written for each paper (N=32), the post-survey, and the interview transcripts. The research team analyzed these combined sources of data to reveal insights.
@@
To measure the quality of the review, we recruited two experts who have conducted research for more than three years of review experience in HCI or CSCW conferences. They counted the number of strengths and weakness in the review (a proxy for coverage) and rated the quality of all final reviews (N=32) with a five-dimension rubric based on reviewer guidelines for CSCW conference and previous research (Yuan et al., 2021):
@@
Two experts rated these five dimensions on a simple seven-point scale (1-7). Each expert first read one paper and wrote a peer review of the paper. After this process, the research team provided provided instructions and two examples for them to rate and discuss until they reached a consensus on ratings based on the instructions. Then, they rated each dimension of the review on the paper independently. The inter-rater reliability between two experts on all 32 data is moderate where Krippendorff’s alpha is higher than 0.50 (Krippendorff, 2018). The research team then used the average scores by two experts for each dimension.
To construct the proxy of review quality, we further asked the experts to count the number of strengths and weaknesses raised by the participants in each review. Another proxy of the review quality we used is the participants’ self-rated satisfaction with their reviews. After participants submitted the reviews, we asked them to rate their satisfaction with their reviews on a scale of 1 -7 (1 as not satisfied at all and 7 as very satisfied).
@@
To measure participants’ interaction with the tool, we instrumented the interface to collect a range of user activity log data. We collected two timing measures – how long each participant took to finish the review session and how long each participant spent editing the review within the text box. The entire review session time includes reading time and writing time. The ReviewFlow interface also collected interaction data to capture how much each participant interacted with each scaffolding feature, including the times they answered the contextual cues, clicked on the citation summary, checked the citation pop-ups and used the outline summarization feature.
@@
To evaluate users’ preferences for the ReviewFlow experience compared to a baseline plain review editor, we asked participants to fill out a short post-study survey. The survey asked participants to directly compare the perceived usefulness, enjoyment, easiness, and sense of control between the ReviewFlow and baseline system. We collected the level of cognitive demand using NASA TLX on a scale of 1-5 (Hart and Staveland, 1988). We further collected the feeling of control and collaboration on a scale of 1-5. Then we specially asked participants to evaluate each of the scaffolding features. The survey collected their 5-point Likert scale ratings on perceived usefulness, and perceived accuracy for each feature.
Previous research showed that scaffolding can promote learners’ self-efficacy (Yantraprakorn et al., 2013). Here, we measured whether using the scaffold system can improve novice reviewers’ self-efficacy and confidence. We asked each participant to report self-efficacy after using the ReviewFlow system and the Baseline system by answering the question “How confident are you in your ability to write a conference peer review next time after using the system (1 as not confident and 7 as very confident)”.
After the post-survey, we conducted a 15-minute semi-structured interview with all participants to capture their overall thoughts as well as specific perceptions of machine-generated highlights and summaries. For example, the research team asked “What do you think of the difference between the task with and without the support of ReviewFlow”, “What did you learn after writing the review with the system”, and “What concerns did you have when using ReviewFlow?”.
Users’ knowledge of each paper’s topic as control variables
We measured participants’ existing knowledge of the two papers as a control variable. Participants described their knowledge of each paper’s topic from not familiar as 1 to very familiar as 7 in the post-survey. All participants reported that they had never read and remembered the papers before. The average rating of knowledge level on the first paper is 3.46 and the average rating for the second paper is 3.42.